# A Comprehensive Study on Large Language Models for Mutation Testing

### v3

@@ 4.1.4. Performance of Using the Same Number of Mutations
In Section 3, we did not restrict the number of mutations generated by each approach. To study the performance of each approach under fixed quantity conditions, we followed the setting of existing study (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by Bert (Degiovanni and Papadakis, 2022), which is 11397. For counts exceeding this number, we randomly selected the specified number of mutations for analysis. We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 16. Therefore, under fixed quantity conditions, the mutations generated by two GPT-4 LLMs are still the closest to real bugs in behavior.
| Metric | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| Real Bug Det. | 31.6% | 68.8% | 64.6% | 66.0% | 82.2% | 87.4% | 88.6% | 87.0% | 33.9% | 68.5% |
| Coupling Rate | 17.4% | 45.4% | 37.2% | 43.6% | 45.6% | 48.8% | 47.9% | 47.9% | 27.1% | 43.6% |
| Ochiai Coefficient | 12.6% | 22.2% | 14.1% | 19.0% | 26.5% | 36.0% | 31.8% | 34.1% | 15.4% | 16.9% |
| Comp. Rate | — | 96.9% | 54.1% | 26.1% | 61.2% | 75.5% | 74.3% | 74.6% | 12.4% | 69.5% |
| Dup Mut. Rate | — | 0.0% | 2.6% | 1.9% | 9.9% | 7.7% | 7.1% | 7.6% | 9.6% | 38.3% |
@@ 4.2. Implications
Based on our findings, we discuss the pros and cons of traditional and LLM-based mutation generation approaches, and suggest potential improvements.
@@ 4.2.1. Advantages of LLM-based Approaches
Additionally, our study explores how different prompt designs, model choices, and contextual information influence mutation generation performance. We find that selecting models with stronger code generation capabilities (e.g., GPT-4o) and designing prompts carefully are critical to achieving optimal results.
@@ 4.2.3. Challenges and Future Directions
Although LLMs demonstrate superior behavioral similarity to real bugs, our findings indicate substantial room for improvement. One major limitation is the high rate of non-compilable mutations, which limits their practical application in mutation testing. Our analysis reveals that common causes include syntactic errors in generated expressions, incorrect method invocations, and type mismatches. To address these challenges, future research should explore integrating LLM-based approaches with program analysis and repair techniques to refine mutation generation. Specifically, leveraging static analysis to filter or adjust syntactically incorrect mutations before execution can enhance their reliability.
@@ 5. Threats to Validity
The selected programming language, mutation generation approaches, datasets, and LLMs could be a threat to the validity of our results. To mitigate this threat, employ the most widely studied mutation testing approaches as baselines, including learning-based and rule-based, the most widely used language (i.e., Java), and the most widely used dataset Defects4J, and we adopt the widely studied models covering a range of base models (i.e., GPT, StarCoder, CodeLlama, and DeepSeek).
@@
Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014a) may be covered in the training set of the studied LLMs. To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2024) which includes programs and faults that were made after the release time of the LLMs we use and thus have limited data leakage risk. Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults. We hypothesize that in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate. On Defects4J, the results are GPT-3.5 (133), GPT-4o (63), GPT-4o-Mini (138), DeepSeek (66), StarChat (13), CodeLlama (28), LEAM (287), Bert (12), and Major (22), respectively. On ConDefects, the results are GPT-3.5 (22), GPT-4o (25), GPT-4o-Mini (31), DeepSeek (32), StarChat (35), CodeLlama (56), LEAM (56), Bert (5), and Major (92), respectively. The comparison indicates that on Defects4J, GPTs and LEAM tend to produce significantly more exact matches than the other approaches. Interestingly, Major produces a similar number of exact matches with Codellama. Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except from GPTs and LEAM) due to exact matches (in the case of Defects4J). Perhaps more interestingly, on the ConDefects dataset, which has not been seen by any of the tools, Major has the highest number of exact matches, indicating a minor influence of any data leakage on the reported results. Nevertheless, the LLMs we studied exhibit the same trend on the two datasets, achieving the Spearman coefficient of 0.943 and the Pearson correlation of 0.944, both with -value less than 0.05, indicating their performance is similar on the two datasets.
@@
The different experimental settings may also threaten the validity of our results. To address the threat, we explore the impacts of prompts, context length, few-shot examples, and mutation numbers on the performance of LLMs. The results show that different settings are highly similar. One potential threat to validity is that the metrics used in this study may not fully reflect the actual bug detection in practice. Our evaluation is conducted on the fixed versions of programs, rather than the faulty versions where real-world testing would typically be applied, which may introduce potential differences. In other words, our study operates under the clean program assumption, which may not fully capture the dynamics of bug detection in practical scenarios. However, to mitigate this threat, we use the widely adopted Defects4J 2.0 dataset, which consists of real bugs from well-maintained open-source projects, ensuring the relevance of our findings to real bugs,
@@
The subjective nature of human decisions in labeling for equivalent mutations and non-compilation errors is another potential threat. To mitigate this threat, we follow a rigorous annotation process where two co-authors independently annotated each mutation. The final Cohen’s Kappa coefficient indicates a relatively high level of agreement between the two annotators.
@@ 6.1. Mutation Testing and Mutation Generation
Mutation testing is a systematic testing approach that aims to guide the creation of high-utility test suites. It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test. By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978) one can identify weaknesses of their test suites, i.e., mutants that are not killed. Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017). Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014b; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015; Wang et al., 2025), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
@@
All these tasks require high-utility mutations. This is because it is required to compile and execute every mutation with tests, which is computationally expensive. Therefore, generating redundant or non-compilable mutations, as done by most of the existing learning-based approaches, increases the computational cost involved (Zhang et al., 2018; Wang et al., 2017, 2021; Garg et al., 2022). Additionally, when mutations are used as substitutes for real defects, the mutations must be as syntactically close to the real bugs as possible, making the mutations more natural (Hindle et al., 2016; Jimenez et al., 2018; Hariri et al., 2019). Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011), the behavior of the mutations needs to couple with real faults/bugs to allow these techniques to be effective.
@@
Traditional mutation testing approaches generate mutations by pre-defined program transformation rules (i.e., mutation operators), which are referred to rule-based. For example, one widely used mutation operator is to change one arithmetic binary operator with another one (e.g., ). During the last two decades, the majority of existing mutation testing approaches have been rule-based (Jia and Harman, 2010; Amalfitano et al., 2022). Researchers primarily focus on approaches in Java (e.g., PIT (Coles et al., 2016), Major (Just et al., 2011), JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), IBIR (Khanfir et al., 2023), etc.) and C/C++ (e.g., Proteum (Delamaro et al., 1996), Milu (Jia and Harman, 2008), WinMut (Wang et al., 2021), etc.), leveraging the mature ecosystems and tooling chains of these languages for mutation testing.
@@
Recently, with the emergence of machine learning techniques, researchers proposed using deep learning models to generate mutations. For instance, DeepMutation (Tufano et al., 2019) utilizes machine sequence-to-sequence neural translation methods, while LEAM (Tian et al., 2022) employs learning-guided rule-based expansion. However, these approaches suffer from the challenges in generating syntactic correct mutations and mutations with diverse forms and often fail to generate syntactically valid mutations. Perhaps the closest to our work is that of Degiovanni et al. who proposed Bert (Degiovanni and Papadakis, 2022). Bert generates mutation by masking the code elements to be mutated and then predicting the mutation via Bert (Devlin et al., 2018).
@@
Several recent studies have also explored the use of LLMs for mutation generation. Tip et al. concurrently worked on using LLMs for mutation generation for JavaScript, but their approach neither included a comprehensive comparison with existing techniques nor did they utilize only three open-source LLMs (Tip et al., 2024). Similarly, Endres et al. employed large models for mutation generation to aid deduction (Endres et al., 2024), while Deb et al. used GPT as one of the baselines in their approach (DEB et al., 2024). In contrast to their projects, our work systematically explores a variety of LLMs and prompt designs, offering a thorough comparison with state-of-the-art approaches. Dakhel et al.(Dakhel et al., 2024) and Ibrahimzada et al.(Ibrahimzada et al., 2023) also proposed generating artificial bugs using LLMs. These approaches mainly evaluate from a certain perspective, such as mutation score or program repair. Compared with these approaches, our study is more comprehensive and evaluates existing mutation generation approaches from many aspects. Specifically, we measure the behavioral similarity between real bugs and mutations using established metrics, evaluate the syntactic diversity of mutations, and analyze the non-compilable mutations produced by LLMs. Our findings illustrate both the strengths and limitations of LLMs for mutation generation, and highlight directions for future improvement.

### v4

@@ 4.1.4. Performance of Using the Same Number of Mutations
In Section 3, we did not restrict the number of mutations generated by each approach. To study the performance of each approach under fixed quantity conditions, we followed the setting of the existing studies (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by BERT (Degiovanni and Papadakis, 2022), which is 11397. This ensures a fair comparison across approaches by controlling for differences in mutation counts, thereby eliminating the potential bias that methods generating more mutations might appear to perform better simply due to quantity rather than quality. For counts exceeding this number, we randomly selected the specified number of mutations for analysis. We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 18. Therefore, under fixed quantity conditions, the mutations generated by two GPT-4 LLMs are still the closest to real bugs in behavior.
| Name | R. B. Detec. | Ochiai Coeffi. | Coup. Rate | Comp. Rate | Dup. Rate |
| PIT | 31.6% | 12.6% | 17.4% | — | — |
| Major | 68.8% | 22.2% | 45.4% | 96.9% | 0.0% |
| LEAM | 64.6% | 14.1% | 37.2% | 54.1% | 2.6% |
| BERT | 66.0% | 19.0% | 43.6% | 26.1 | 1.9% |
| BugFarm (GPT-4o) | 58.6% | 17.7% | 47.9% | 74.2% | 7.5% |
| LLMorpheus (GPT-4o) | 67.5% | 23.6% | 50.2% | 68.2% | 7.3% |
| LLMut (GPT-3.5) | 82.2% | 26.5% | 45.6% | 61.2% | 9.9% |
| LLMut (GPT-4o) | 87.4% | 36.0% | 48.8% | 75.5% | 7.7% |
| LLMut (GPT-4o-M) | 88.6% | 31.8% | 47.9% | 74.3% | 7.1% |
| LLMut (DS-761b) | 89.2% | 39.4% | 49.4% | 76.3% | 7.7% |
| LLMut (DS-236b) | 87.0% | 34.1% | 47.9% | 74.6% | 7.6% |
| LLMut (SC-16b) | 33.9% | 15.4% | 27.1% | 12.4% | 9.6% |
| LLMut (CL-13b) | 68.5% | 16.9% | 43.6% | 69.5% | 38.3% |
| Metric | LLMut (GPT-3.5) | LLMut (DS-761b) | ||||
| Default | More Detailed | Default + 10 Err. | Default | More Detailed | Default + 10 Err. | |
| Real Bug Detec | 0.917 | 0.900 | 0.900 | 0.950 | 0.950 | 0.933 |
| Avg. Ochiai | 0.467 | 0.483 | 0.467 | 0.500 | 0.500 | 0.517 |
| Coupling Rate | 0.136 | 0.129 | 0.117 | 16.9% | 17.8% | 16.0% |
| Comp. Rate | 67.5% | 71.1% | 68.7% | 78.4% | 76.4% | 77.5% |
| Dup.Mut.Rate | 13.7% | 14.1% | 10.4% | 5.1% | 4.1% | 6.2% |
Note: The prompt achieving the best performance in terms of a metric is highlighted in grey background color.
4.2. Impact of Prompt Turbulence on LLM-Based Mutation Generation
In software engineering practice, even when using the same information sources, users may design prompts differently, for example, varying the level of detail in the requirements description or introducing minor language errors. To investigate the impact of such variations, we derived two prompt variants from the default prompt shown in Figure 1.
More Detailed Output Specification: replacing the output specifications of the default prompt by explicitly requiring: Prohibit generating identical mutants, i.e., ensure ‘aftercode’ ‘precode’, and all mutants are unique. Write all output mutants in a JSON file.
Default + 10 Errors: randomly injecting 10 English grammar mistakes into the default prompt, such as subject–verb agreement errors (e.g., it are) and noun number errors (e.g., two mutation).
Following the same experimental setup as before, we compared these two variants using GPT-3.5 and DeepSeek-V3-761b, with the results summarized in Table 19.
For GPT-3.5, the More Detailed Output Specification variant slightly improves the compilation rate (+3.6%) but also increases the duplicate mutation rate. In contrast, the Default + 10 Errors variant reduces real bug detection and coupling rate compared to the default prompt. Interestingly, the Default + 10 Errors variant, presumably the lowest in prompt quality, achieves the best performance on duplicate mutation rate. For DeepSeek-V3-761b, the differences across variants are minimal. For instance, the default and the More Detailed Output Specification prompts achieve identical real bug detection and average Ochiai, and the More Detailed Output Specification variant even shows a small increase in average Ochiai.
Overall, these results suggest that prompt wording has a certain degree of randomness in its effect on mutation generation performance. To further verify this, we conducted similarity analysis across the three prompts for each metric and found no statistically significant superiority of any one prompt over the other two.
| Metric | LLMut (GPT-3.5) | LLMut (GPT-4o) | LLMut (DS-671b) | |||
| FOM | HOM | FOM | HOM | FOM | HOM | |
| Real Bug Detec. | 0.917 | 0.900 | 0.933 | 0.933 | 0.950 | 0.933 |
| Avg. Ochiai | 0.467 | 0.483 | 0.467 | 0.483 | 0.500 | 0.483 |
| Coupling Rate | 0.136 | 0.249 | 0.151 | 0.307 | 0.169 | 0.306 |
| Compilability Rate | 67.5% | 63.4% | 76.3% | 70.2% | 78.4% | 73.3% |
| Dup. Mut. Rate | 13.7% | 5.2% | 5.3% | 3.5% | 5.1% | 3.1% |
Note: The order achieving the best performance in terms of a metric is highlighted in grey background color.
4.3. Effectiveness on High-Order Mutation Generation
Previously, we only evaluated LLMs in generating first-order mutations (FOM). To explore the effectiveness of LLMs in generating high-order mutations (HOMs) (Jia and Harman, 2009; Wong et al., 2020), we conducted a preliminary experiment by evaluating the recent mainstream models, including GPT-3.5, GPT-4o, and DeepSeek-V3-761b.
We modified the original prompt in Figure 1 to explicitly request HOM generation. Table 20 shows the results. For the effectiveness metrics, HOMs generally achieve average Ochiai scores and higher coupling rates across these models, indicating closer semantic alignment with real bugs. However, FOMs achieve higher real bug detectability on all models. For the validity metrics, HOMs exhibit lower duplicate mutation rates, suggesting improved mutation diversity. However, FOMs exhibit significantly higher compilability rates, which aligns with the expectation that larger code changes are more prone to introducing compilation errors. Our preliminary experiments suggest that strengthening LLM-based HOM generation is promising, as HOMs are semantically closer to real bugs and face more challenges in maintaining syntactic correctness.
@@ 4.4. Implications
Based on our findings, we discuss the pros and cons of traditional and LLM-based mutation generation approaches and suggest potential improvements.
@@ 4.4.1. Advantages of LLM-based Approaches
Additionally, our study explores how different prompt designs, prompt turbulence, model choices, and contextual information influence mutation generation performance. We find that selecting models with stronger code generation capabilities (e.g., GPT-4o) and designing prompts carefully are critical to achieving optimal results.
Moreover, our study demonstrates that LLMs have clear potential in generating HOMs. Compared with first-order mutations (FOMs), HOMs generally achieve higher coupling rates and average Ochiai scores, suggesting that they are semantically closer to real bugs, while suffering from lower compilability rates and reduced real bug detectability. The results highlight that strengthening LLM-based HOM generation is therefore a promising research direction.
@@ 4.4.3. Challenges and Future Directions
Although LLMs demonstrate superior behavioral similarity to real bugs, our findings indicate that there is still substantial room for improvement. One major limitation is the high rate of non-compilable mutations, which limits their practical application in mutation testing. Our analysis reveals that common causes include syntactic errors in generated expressions, incorrect method invocations, and type mismatches. To address these challenges, future research should explore the integration of LLM-based approaches with program analysis and repair techniques to refine mutation generation. Specifically, leveraging static analysis to filter or adjust syntactically incorrect mutations before execution can enhance their reliability.
Moreover, exploring the applications of LLM-generated mutation in the downstream tasks, such as mutation-based test case prioritization (MBTCP) and mutation-based fault localization (MBFL). Our study has shown that LLM-generated mutations often have a higher similarity with real bugs (e.g., have larger Ochiai coefficients), highlighting the potential of adopting them for these tasks.
Another important direction is cross-language generalization. While our study primarily focuses on Java, mutation testing is equally critical in languages such as C/C++, Python, JavaScript, and Rust. Traditional mutation testing tools usually require language-specific implementations, leading to duplicated engineering efforts and limited portability. In contrast, modern LLMs possess inherent multilingual capabilities, as they are pre-trained on large-scale, cross-language code corpora. This enables them to naturally transfer mutation strategies across programming languages without retraining from scratch. Future work should investigate how well LLMs can exploit this cross-language knowledge, and how prompts, training data, or filtering strategies can be adapted to different language paradigms (e.g., strongly typed vs. dynamically typed). Advancing towards multilingual mutation generation would not only broaden the applicability of LLMs in software testing but also provide stronger evidence for their robustness in heterogeneous, real-world development environments.
4.5. Limitations of Our Study
Our study has the following limitations that are left for further exploration.
First, while we evaluated a comprehensive set of static and behavioral metrics, we did not investigate more advanced dynamic indicators. For instance, the phenomenon of subsuming mutations (Garg et al., 2022; Diniz et al., 2021), where one mutation is always killed by any test that kills another, has been studied in the mutation testing literature as a way to assess better redundancy and the true fault-detection power of test suites. Incorporating such dynamic analyses would provide deeper insights into the effectiveness of LLM-generated mutations beyond the metrics reported in this paper. Additionally, while in the context of test suite enhancement, two mutants with identical failing test sets may be considered redundant, in mutation-based fault localization (Wang et al., 2025), the number of mutants per statement directly influences suspiciousness scores, since these are typically computed as averages across all associated mutants.
Second, our experiments are limited to the Java programming language. Although Java is the most widely studied language in mutation testing, it does not cover the diversity of programming paradigms. Mutation behaviors may differ significantly in languages such as C/C++, Python, JavaScript, and Rust, due to differences in type systems, memory models, and runtime environments. Future research should examine the generalizability of LLM-based mutation generation across multiple languages. Notably, modern LLMs are inherently multilingual, trained on large-scale cross-language corpora, which makes cross-language mutation generation a promising direction for broadening the applicability of our findings.
Third, our study relies on fixed few-shot examples, which restricts LLMs to producing context-specific mutations for the given code snippets. In addition, we impose no restrictions on the mutation operators applied by LLMs, whereas rule-based approaches rely on predefined and relatively simple operators. Consequently, it is still an open question how LLMs would perform if constrained to the same operator set as rule-based approaches.
@@ 5. Threats to Validity
The selected programming language, mutation generation approaches, datasets, and LLMs could be a threat to the validity of our results. To mitigate this threat, employ the most widely studied mutation testing approaches as baselines, including learning-based and rule-based, the most widely used language (i.e., Java), and the most widely used dataset Defects4J, and we adopt the widely studied models covering a range of base models (i.e., GPT, StarCoder, CodeLlama, and DeepSeek). Moreover, our study further adopts two state-of-the-art LLM-based mutation generation approaches (i.e., BugFarm and LLMorpheus) as baselines.
The metrics we used may threaten the conclusion of our study. Each metric has inherent limitations. For instance, Real Bug Detection Rate and Coupling Rate capture only partial aspects of similarity to real bugs: a method may detect many bugs but with weak coupling on each bug, or achieve strong coupling concentrated on a few bugs. Likewise, AST Edit Distance reflects structural differences, but a higher syntactic distance does not necessarily imply larger semantic changes. Additionally, the measurement of Average Generation Time may be affected by external factors such as GPU type, service provider implementation, and network fluctuations, introducing unavoidable noise. To mitigate the bias of individual metrics, we report a comprehensive set of complementary measures. Particularly, besides time cost, we also report token consumption, which provides a more stable and model-independent estimate of computational and financial cost.
@@
Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014b) may be covered in the training set of the studied LLMs. To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2024), which includes programs and faults that were made after the release time of the most LLMs we use (expect DeepSeek-V3-761b only) and thus have limited data leakage risk. Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults. We hypothesize that, in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate. On Defects4J, the results are GPT-3.5 (133), GPT-4o (63), GPT-4o-Mini (138), DeepSeek-V3-761b (72), DeepSeek-V2-236b (66), BugFarm (59), LLMorpheus (26), StarChat (13), CodeLlama (28), LEAM (287), Bert (12), and Major (22), respectively. On ConDefects, the results are GPT-3.5 (22), GPT-4o (25), GPT-4o-Mini (31), DeepSeek-V3-761b (35), DeepSeek-V2-236b (32), BugFarm (22), LLMorpheus (19), StarChat (35), CodeLlama (56), LEAM (56), Bert (5), and Major (92), respectively. The comparison indicates that on Defects4J, GPTs and LEAM tend to produce significantly more exact matches than the other approaches. Interestingly, Major produces a similar number of exact matches with Codellama. Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except for GPTs and LEAM) due to exact matches (in the case of Defects4J). Perhaps more interestingly, on the ConDefects dataset, which has not been seen by any of the tools, Major has the highest number of exact matches, indicating a minor influence of any data leakage on the reported results. Nevertheless, the LLMs we studied exhibit the same trend on the two datasets, achieving the Spearman coefficient of 0.943 and the Pearson correlation of 0.944, both with -value less than 0.05, indicating their performance is similar on the two datasets.
@@
The different experimental settings may also threaten the validity of our results. To address the threat, we explore the impacts of prompt design, prompt turbulence, context length, few-shot examples, and mutation numbers on the performance of LLMs. The results show that different settings are highly similar. One potential threat to validity is that the metrics used in this study may not fully reflect the actual bug detection in practice. Our evaluation is conducted on the fixed versions of programs, rather than the faulty versions where real-world testing would typically be applied, which may introduce potential differences. In other words, our study operates under the clean program assumption, which may not fully capture the dynamics of bug detection in practical scenarios. However, to mitigate this threat, we use the widely adopted Defects4J 2.0 dataset, which consists of real bugs from well-maintained open-source projects, ensuring the relevance of our findings to real bugs.
@@
The subjective nature of human decisions in labeling for equivalent mutations, non-compilation errors, and alive mutations is another potential threat. To mitigate this threat, we follow a rigorous annotation process where two co-authors independently annotated each mutation. The final Cohen’s Kappa coefficient indicates a relatively high level of agreement between the two annotators.
@@ 6.1. Mutation Testing and Mutation Generation
Mutation testing is a systematic testing approach that aims to guide the creation of high-utility test suites. It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test. By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978), one can identify weaknesses of their test suites, i.e., mutants that are not killed. Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017). Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014c; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015; Wang et al., 2025), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
@@
All these tasks require high-utility mutations. This is because it is required to compile and execute every mutation with tests, which is computationally expensive. Therefore, generating redundant or non-compilable mutations, as done by most of the existing learning-based approaches, increases the computational cost involved (Zhang et al., 2018; Wang et al., 2017, 2021; Garg et al., 2022). Additionally, when mutations are used as substitutes for real defects, the mutations must be as syntactically close to the real bugs as possible, making the mutations more natural (Hindle et al., 2016; Jimenez et al., 2018; Hariri et al., 2019). Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011), the behavior of the mutations needs to be coupled with real faults/bugs to allow these techniques to be effective.
@@
Traditional mutation testing approaches generate mutations by pre-defined program transformation rules (i.e., mutation operators), which are referred to as rule-based. For example, one widely used mutation operator is to change one arithmetic binary operator with another one (e.g., ). During the last two decades, the majority of existing mutation testing approaches have been rule-based (Jia and Harman, 2010; Amalfitano et al., 2022). Researchers primarily focus on approaches in Java (e.g., PIT (Coles et al., 2016), Major (Just et al., 2011), JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), IBIR (Khanfir et al., 2023), etc.) and C/C++ (e.g., Proteum (Delamaro et al., 1996), Milu (Jia and Harman, 2008), WinMut (Wang et al., 2021), etc.), leveraging the mature ecosystems and tooling chains of these languages for mutation testing.
@@
Recently, with the emergence of machine learning techniques, researchers proposed using deep learning models to generate mutations. For instance, DeepMutation (Tufano et al., 2019) utilizes machine sequence-to-sequence neural translation methods, while LEAM (Tian et al., 2022) employs learning-guided rule-based expansion. LEAM++, the extension of LEAM, further explores learning-based mutation selection (Tian et al., [n. d.]). However, these approaches suffer from the challenges in generating syntactically correct mutations and mutations with diverse forms and often fail to generate syntactically valid mutations. Perhaps the closest to our work is that of Degiovanni et al., who proposed BERT (Degiovanni and Papadakis, 2022). BERT generates mutation by masking the code elements to be mutated and then predicting the mutation via BERT (Devlin et al., 2018).
@@
Several recent studies have also explored the use of LLMs for mutation generation. Tip et al. concurrently investigated the use of LLMs for mutation generation in JavaScript (Tip et al., 2025), primarily focusing on the feasibility of leveraging LLMs to produce mutations. In contrast, our study conducts a broader comparison by evaluating LLMs against a diverse set of baselines, including both rule-based and learning-based approaches. Furthermore, we explore different prompt strategies, such as incorporating few-shot examples and enforcing structured JSON output, and we systematically examine multiple prompt templates in the context of mutation testing. Similarly, Endres et al. employed large models for mutation generation to aid deduction (Endres et al., 2024), while Deb et al. used GPT as one of the baselines in their approach (DEB et al., 2024). In contrast to their projects, our work systematically explores a variety of LLMs and prompt designs, offering a thorough comparison with state-of-the-art approaches. Dakhel et al. (Dakhel et al., 2024) and Ibrahimzada et al. (Ibrahimzada et al., 2023) also proposed generating artificial bugs using LLMs. These approaches mainly evaluate from a certain perspective, such as mutation score or program repair. BugFarm (Ibrahimzada et al., 2023) intends to generate artificial bugs that are hard to kill, by first choosing code locations based on a trained small model and then mutating the code via an LLM. Compared with these approaches, our study is more comprehensive and evaluates existing mutation generation approaches from many aspects. Specifically, we measure the behavioral similarity between real bugs and mutations using established metrics, evaluate the syntactic diversity of mutations, and analyze the non-compilable mutations produced by LLMs. Our findings illustrate both the strengths and limitations of LLMs for mutation generation, and highlight directions for future improvement.